OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2

OpenStudy (amorfide):

you asked htis question i answered http://openstudy.com/users/angelsundoing12#/updates/4fba81f7e4b05565342f70f4

OpenStudy (anonymous):

i tried it with the y2-y1 got a little frustrated

OpenStudy (anonymous):

figured id ask again

OpenStudy (amorfide):

what do you not understand? let me know so i can explain

OpenStudy (anonymous):

i got 14/-7 thats where im stuck

OpenStudy (amorfide):

that cancels down... to -2 so your gradient is -2 do you agree?

OpenStudy (anonymous):

yes

OpenStudy (amorfide):

okay so now choose one set of coordinates lets use (2, −2) y-y1=m(x-x1) y1= y coordinate x1=x coordinate m=gradient sub in your values

OpenStudy (apoorvk):

a line passing through two points (x',y') and (x'',y'') has the formula \[\frac{y-y'}{x-x'} = \frac{y'-y''}{y'-y''}\]

OpenStudy (amorfide):

oh i assume im wrong? oops

OpenStudy (anonymous):

i giv up im just gonna guess :(